# PromptKD: Distilling Student-Friendly Knowledge for Generative Language Models via Prompt Tuning

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have raised concerns about inference costs, increasing the need for research into model compression. While knowledge distillation (KD) is a prominent method for this, research on KD for generative language models like LLMs is relatively sparse, and the approach of distilling student-friendly knowledge, which has shown promising performance in KD for classification models, remains unexplored in generative language models. To explore this approach, we propose PromptKD, a simple yet effective method that utilizes prompt tuning - for the first time in KD - to enable generative language models to transfer student-friendly knowledge. Unlike previous works in classification that require fine-tuning the entire teacher model for extracting student-friendly knowledge, PromptKD achieves similar effects by adding a small number of prompt tokens and tuning only the prompt with student guidance. Extensive experiments on instruction-following datasets using the GPT-2 model family show that PromptKD achieves state-of-the-art performance while adding only 0.0007% of the teacher’s parameters as prompts. Further analysis suggests that distilling student-friendly knowledge alleviates exposure bias effectively throughout the entire training process, leading to performance enhancements.
@@ 1 Introduction
With the massive improvement of generative language models, such as the emerging abilities (Wei et al., 2022) observed in large language models (LLMs), there is a growing need for model compression research to efficiently deploy models in various tasks (Touvron et al., 2023; Taori et al., 2023). However, among notable compression methods such as knowledge distillation (KD; Hinton et al., 2015; Kim and Rush, 2016; Gu et al., 2024), pruning (Ma et al., 2023), and quantization (Tao et al., 2022), KD has not been successfully applied to generative language models.
@@
For evaluation, we measure the instruction-following performance (Ouyang et al., 2022), aiming to cover a variety of tasks that generative language models can perform. Compared to the existing baseline, PromptKD achieves state-of-the-art performance by adding prompt parameters equivalent to only 0.0007% of the teacher parameters, as depicted in Figure 1. Additionally, the analysis of exposure bias suggests that remarkable alleviation of exposure bias through student-friendly knowledge is likely the cause of performance improvement. Lastly, through ablation studies, we confirm the necessity of regularization loss and the importance of prompt initialization.
@@ 2 Related Work
For text generation, Sanh et al. (2019) minimizes the KL divergence between the next token prediction distributions of the teacher and student at each time step. In addition, some research (Calderon et al., 2023; Agarwal et al., 2024) focus on the sentences inputted to the teacher and student during the distillation process. For example, Kim and Rush (2016) uses sentences generated by the teacher as pseudo-targets instead of ground truth. Moreover, black-box KD methods (Hsieh et al., 2023; Ho et al., 2023) that use inference-only black-box LLMs as teachers and augment existing data before training are proposed. Recently, Agarwal et al. (2024); Gu et al. (2024) explored discrepancy metrics between model distributions and used sentences generated by the student as pseudo-targets to minimize exposure bias. However, there have been no attempts yet to distill student-friendly knowledge while the teacher is aware of the student’s capacity. Although Liang et al. (2023b) incorporates task-aware filters into both teacher and student to transfer knowledge, its scalability is limited due to the addition of filters at each layer for layer distillation. Crucially, it encourages knowledge to be task-specific, making it diverges from what we aim to explore in this paper.
@@ 3.2 Prompt Tuning for Adaptive Teaching
#Params Method Instruction-following datasets
@@
MiniLLM 24.2 12.7 ${}^{\dagger}$ 25.1 26.2
GKD 24.8 11.1 ${}^{\dagger}$ 20.7 23.2
PromptKD 25.6 13.1 ${}^{\dagger}$ 26.8 28.9
@@
MiniLLM 26.3 ${}^{\dagger}$ ${}^{\dagger}$ 26.4 31.2
GKD 26.9 ${}^{\dagger}$ ${}^{\dagger}$ 26.6 30.9
PromptKD ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
MiniLLM 26.2 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
GKD 26.9 14.1 ${}^{\dagger}$ 25.4 29.6
PromptKD 26.9 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
Table 1: Evaluation results on 5 instruction-following datasets. Each ROUGE-L score is averaged over 5 random seeds. The best score for each model size is highlighted in boldface. ${}^{\dagger}$esults surpass those of the teacher.
@@ 4.1 Experimental Setup
We split the Dolly (Conover et al., 2023), consisting of 15,000 human-written instruction-response pairs, into 14,000 for training and 500 for validation and testing. For evaluation, we employ not only the Dolly but also 4 additional datasets: SelfInst (Wang et al., 2023), consisting of user-oriented instruction-following sets; Vicuna (Chiang et al., 2023), comprising 80 questions used in the Vicuna evaluation; S-NI, the test set of Super-NaturalInstructions (Wang et al., 2022); and UnNI, the core dataset of UnnaturalInstructions (Honovich et al., 2023). Similar to Gu et al. (2024), data samples with ground truth response lengths of 11 or more are utilized for S-NI and UnNI. We generate 5 responses for each request in each dataset using different random seeds and evaluate them to report the average scores for reliability. We choose the ROUGE-L score Lin (2004) as the metric for evaluation, as it aligns well with human preferences (Wang et al., 2022) in instruction-following evaluations. The best checkpoint based on the ROUGE-L score on the validation set is used for evaluation.
@@
Models
To evaluate the instruction-following performance of PromptKD, we utilize pre-trained models from the GPT-2 family. GPT-2 XL (1.5B params) is employed for the teacher model, and GPT-2 Base (120M params), GPT-2 Medium (340M params), GPT-2 Large (760M params) are used for the student model. Before knowledge distillation, the teacher model undergoes supervised fine-tuning on the Dolly training set. Similarly, the student model is also fine-tuned on the same training data for only three epochs, following the previous works (Agarwal et al., 2024; Gu et al., 2024).
Baselines
PromptKD is compared with various approaches ranging from supervised fine-tuning (SFT), which does not involve knowledge distillation, to commonly used methods in generation tasks such as Supervised KD (KD; Sanh et al., 2019), SeqKD (Kim and Rush, 2016), and more recent proposals like MiniLLM (Gu et al., 2024) and GKD (Agarwal et al., 2024). KD and SeqKD both aim to minimize the discrepancy between the model distributions of teacher and student at each token step. The difference lies in whether the input sentence is ground truth or pseudo-target generated by the teacher. MiniLLM replaces forward KL divergence with reverse KL divergence and updates the student model using policy gradient. On the other hand, GKD focuses on distribution discrepancy metrics and pseudo-targets to propose a general method. In this paper, GKD computes reverse KL divergence and utilizes sentences generated by the student as pseudo-targets, and this choice is based on the reported performance in their paper. Additionally, it is worth noting that the students for MiniLLM, GKD, and PromptKD all commence from the same supervised fine-tuned checkpoint, while other methods start from pre-trained models. For training details, please see the Appendix A.
4.2 Experimental Results
Firstly, PromptKD achieves state-of-the-art performance overall in the instruction-following setting, outperforming other KD baselines. Additionally, it also outperforms on 4 datasets not used in training, demonstrating PromptKD’s superb generalization ability. It’s worth noting that despite MiniLLM incorporating language modeling loss through the corpus used for pre-training, PromptKD exhibits better performance.
Lastly, as the model size increases, PromptKD outperforms in more datasets. This can be attributed to the fact that prompt tuning exhibits a similar effect to full-parameter fine-tuning as the model size grows (Lester et al., 2021). Thanks to the scalability and efficiency of prompt tuning, PromptKD can be expected to yield outstanding results even when applied to larger models.
@@ 4.3 Analysis
SFT 2.36 2.75 0.43
KD 5.91 6.45 0.85
SeqKD 5.91 6.45 0.86
MiniLLM 12.63 21.45 7.59
GKD 6.23 6.93 8.48
PromptKD 13.47 14.40 9.78
Table 2: Comparison of computational costs. Where MA denotes the maximum allocated memory on the GPU and CA denotes the maximum cached memory on the GPU. Time indicates the total training time for each method. All computational costs are calculated on 8 NVIDIA GeForce RTX 3090 (24 GB) GPUs.
@@
To demonstrate the efficiency of PromptKD, we compare its computational cost with baselines in Table 2. GPT-2 XL and GPT-2 Base are used as the teacher and the student, with measurements conducted on 8 NVIDIA GeForce RTX 3090 GPUs. From a time perspective, methods that sample the student at each iteration to create pseudo-targets take significantly more time than those that do not. Additionally, while PromptKD introduces only a small amount of additional parameters, namely the product of the prompt length and input embedding dimension, it requires a considerable amount of memory due to back-propagation during training. In contrast, MiniLLM, which does not add parameters, requires more cached memory than PromptKD. This is because it receives guidance not only from the teacher but also from the corpus used in pre-training and calculates rewards in advance at every iteration. When compared to MiniLLM and GKD, which show superior performance to other baselines, PromptKD shows competitive advantages because it performs reliably better despite requiring similar or slightly more time and memory resources.
4.4 Ablation Study
Regularization loss
#Params w/o $\mathcal{L}_{\text{reg}}$ w/ $\mathcal{L}_{\text{reg}}$
120M 21.97 22.25
340M 24.13 23.92
760M 24.47 25.08
Prompt settings
Although the regularization loss effectively mitigates the initial instability, the prompt’s length and initialization also significantly influence the prompt tuning process (Hou et al., 2022). Therefore, the average instruction-following performance is measured by varying the prompt length $m$ from 5, 7, 10 and the initialization method from random, padding, text. Results are summarized in Figure 4. In the padding method, all prompt tokens are initialized with the embedding of the "[PAD]" token, while in the text method, the sentence "Suppose you are a student." is tokenized, and these embeddings are used for initializing prompt tokens from the beginning. In this case, if the number of prompt tokens is smaller, the sentence is truncated, while if it is larger, all embeddings of the sentence are assigned, and then the embeddings are assigned again from the beginning for the next prompt token.
KL divergences
To assess the impact of distribution discrepancy metrics, we conduct ablation study on this. We observe the best performance when measuring discrepancy using reverse KL divergence in both KD loss $\mathcal{L}_{\text{kd}}$ and regularization loss $\mathcal{L}_{\text{reg}}$, similar to previous observations. Detailed experimental results are provided in the Appendix C.
@@ 5 Conclusions
In this work, we have pioneered the exploration of extracting and distilling student-friendly knowledge for generative language models. To achieve this, we have proposed a novel method called PromptKD, which leverages prompt tuning in knowledge distillation for the first time. Thanks to the memory-efficient nature of prompts and the advantage of replacing full-parameter fine-tuning, particularly beneficial for larger models like LLMs, PromptKD has proven to be an efficient approach. Through extensive experiments, PromptKD has achieved state-of-the-art performance, confirming the effectiveness of student-friendly knowledge in generation tasks. Additionally, through exposure bias analysis, we have demonstrated that PromptKD successfully alleviates exposure bias throughout the training process.
@@ Appendix A Training Details
In our study, we employ the AdamW (Loshchilov and Hutter, 2019) optimizer for training, with batch sizes of 32 for GPT-2 Base and 8 for GPT-2 Medium and Large. The learning rates of prompt and student are set at 5e-5 for Base, 1e-5 for Medium, and 5e-6 for Large. For the generation, we sample with top-k and top-p parameters at 0 and 1.0, respectively, and use a fixed temperature of 1.0. Training and generation phases both have a maximum sequence length of 512 and a maximum prompt length of 256. Please note that we pre-process each instruction-following dataset by converting the instruction-response pairs into a standardized sentence structure, as shown in Table 4. For reproducibility of our PromptKD, we will make both the code and the checkpoints public.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have raised concerns about inference costs, increasing the need for research into model compression. While knowledge distillation (KD) is a prominent method for this, research on KD for generative language models like LLMs is relatively sparse, and the approach of distilling student-friendly knowledge, which has shown promising performance in KD for classification models, remains unexplored in generative language models. To explore this approach, we propose PromptKD, a simple yet effective method that utilizes prompt tuning - for the first time in KD - to enable generative language models to transfer student-friendly knowledge. Unlike previous works in classification that require fine-tuning the entire teacher model for extracting student-friendly knowledge, PromptKD achieves similar effects by adding a small number of prompt tokens and tuning only the prompt with student guidance. Extensive experiments on instruction-following datasets show that PromptKD achieves state-of-the-art performance while adding only 0.0007% of the teacher’s parameters as prompts. Further analysis suggests that distilling student-friendly knowledge alleviates exposure bias effectively throughout the entire training process, leading to performance enhancements.
@@ 1 Introduction
With the massive improvement of generative language models, such as the emerging abilities (Wei et al., 2022) observed in large language models (LLMs), there is a growing need for model compression research to efficiently deploy models in various tasks (Touvron et al., 2023b; Taori et al., 2023). However, among notable compression methods such as knowledge distillation (KD; Hinton et al., 2015; Kim and Rush, 2016; Gu et al., 2024), pruning (Ma et al., 2023), and quantization (Tao et al., 2022), KD has not been successfully applied to generative language models.
@@
For evaluation, we measure the instruction-following performance (Ouyang et al., 2022), aiming to cover a variety of tasks that generative language models can perform. Compared to the existing baseline, PromptKD achieves state-of-the-art performance by adding prompt parameters equivalent to only 0.0007% of the teacher parameters, as depicted in Figure 1. Additionally, the analysis of exposure bias suggests that remarkable alleviation of exposure bias through student-friendly knowledge is likely the cause of performance improvement. Lastly, we explore the student-friendly knowledge in PromptKD and confirm the necessity of regularization loss and the importance of prompt initialization through ablation studies.
@@ 2 Related Work
For text generation, Sanh et al. (2019) minimizes the KL divergence between the next token prediction distributions of the teacher and student at each time step. In addition, some research (Calderon et al., 2023; Agarwal et al., 2024) focus on the sentences inputted to the teacher and student during the distillation process. For example, Kim and Rush (2016) uses sentences generated by the teacher as pseudo-targets instead of ground truth. Moreover, black-box KD methods (Hsieh et al., 2023; Ho et al., 2023) that use inference-only black-box LLMs as teachers and augment existing data before training are proposed. Recently, Agarwal et al. (2024); Gu et al. (2024) explored discrepancy metrics between model distributions and used sentences generated by the student as pseudo-targets to minimize exposure bias. However, there have been no attempts yet to distill student-friendly knowledge while the teacher is aware of the student’s capacity. Although Liang et al. (2023b) incorporates task-aware filters into both teacher and student to transfer knowledge, its scalability is limited due to the addition of filters at each layer for layer distillation. Crucially, it encourages knowledge to be task-specific, making it diverge from what we aim to explore in this paper.
@@ 3.2 Prompt Tuning for Adaptive Teaching
Model #Params Method Instruction-following datasets
@@
GKD 24.8 11.1 17.7 20.7 23.2
MiniLLM 24.2 12.7 16.9 25.1 26.2
PromptKD 25.6 13.1 16.8 26.8 28.9
@@
GKD 26.9 14.8 17.8 26.6 30.9
MiniLLM 26.3 14.8 17.9 26.4 31.2
PromptKD 27.3 15.0 17.6 27.1 32.6
@@
GKD 26.9 14.1 17.1 25.4 29.6
MiniLLM 26.2 15.8 16.9 28.5 33.5
PromptKD 26.9 16.4 17.8 29.5 34.8
OPT 13B Teacher 29.3 17.7 17.3 30.7 33.8
1.3B MiniLLM 26.8 15.2 18.1 28.6 30.9
PromptKD 28.0 15.5 18.5 29.6 33.5
2.7B MiniLLM 27.2 16.2 18.6 29.8 33.1
PromptKD 28.7 17.8 18.9 31.4 34.8
6.7B MiniLLM 28.6 18.0 19.1 32.5 34.5
PromptKD 29.9 19.0 19.8 33.8 35.2
Llama 13B Teacher 30.2 23.1 19.0 35.7 36.9
7B MiniLLM 29.0 21.3 20.6 36.7 38.1
PromptKD 30.0 23.4 21.1 36.6 38.9
Table 1: Evaluation results on 5 instruction-following datasets. Each ROUGE-L score is averaged over 5 random seeds. The best score for each model size is highlighted in boldface. Results surpass those of the teacher.
@@ 4.1 Experimental Setup
We split the Dolly (Conover et al., 2023), consisting of 15,000 human-written instruction-response pairs, into 14,000 for training and 500 for validation and testing. For evaluation, we employ not only the Dolly but also 4 additional datasets: SelfInst (Wang et al., 2023), consisting of user-oriented instruction-following sets; Vicuna (Chiang et al., 2023), comprising 80 questions used in the Vicuna evaluation; S-NI, the test set of Super-NaturalInstructions (Wang et al., 2022); and UnNI, the core dataset of UnnaturalInstructions (Honovich et al., 2023). Similar to Gu et al. (2024), data samples with ground truth response lengths of 11 or more are utilized for S-NI and UnNI. We generate 5 responses for each request in each dataset using different random seeds and evaluate them to report the average scores for reliability. We choose the ROUGE-L score Lin (2004) as the metric for evaluation, as it aligns well with human preferences (Wang et al., 2022) in instruction-following evaluations. The best checkpoint based on the ROUGE-L score on the validation set is used for evaluation. We also measure the GPT-4 feedback scores (Zheng et al., 2024), which are separately summarized in Appendix C.
Models
To evaluate the instruction-following performance of PromptKD across various models, we utilize pre-trained GPT-2 (Radford et al., 2019), OPT (Zhang et al., 2022), and Llama (Touvron et al., 2023a) model families. For the GPT-2 model family, GPT-2 XL (1.5B params) is employed for the teacher model, and GPT-2 Base (120M params), GPT-2 Medium (340M params), GPT-2 Large (760M params) are used for the student model. For the OPT and Llama model families, we use OPT-13B and Llama-13B as the teacher models, and OPT-1.3B, OPT-2.7B, OPT-6.7B, and Llama-7B as the student models, respectively. Before knowledge distillation, the teacher model undergoes supervised fine-tuning on the Dolly training set. Similarly, the student model is also fine-tuned on the same training data for only three epochs, following the previous works (Agarwal et al., 2024; Gu et al., 2024).
Baselines
PromptKD is compared with various approaches ranging from supervised fine-tuning (SFT), which does not involve knowledge distillation, to commonly used methods in generation tasks such as Supervised KD (KD; Sanh et al., 2019), SeqKD (Kim and Rush, 2016), and more recent proposals like MiniLLM (Gu et al., 2024) and GKD (Agarwal et al., 2024). KD and SeqKD both aim to minimize the discrepancy between the model distributions of teacher and student at each token step. The difference lies in whether the input sentence is ground truth or pseudo-target generated by the teacher. MiniLLM replaces forward KL divergence with reverse KL divergence and updates the student model using policy gradient. On the other hand, GKD focuses on distribution discrepancy metrics and pseudo-targets to propose a general method. In this paper, GKD computes reverse KL divergence and utilizes sentences generated by the student as pseudo-targets, and this choice is based on the reported performance in their paper. Additionally, it is worth noting that the students for MiniLLM, GKD, and PromptKD all commence from the same supervised fine-tuned checkpoint, while other methods start from pre-trained models. Due to resource limitations, experiments on the OPT and Llama models are conducted only in comparison with MiniLLM, which demonstrated outstanding performance among all baselines in the GPT-2 results. For training details, please see the Appendix A.
4.2 Experimental Results
Firstly, PromptKD achieves state-of-the-art performance overall in the instruction-following setting, outperforming other KD baselines. Additionally, it also outperforms on 4 datasets not used in training, demonstrating PromptKD’s superb generalization ability. These results robustly demonstrate the superiority of PromptKD, as they consistently appear across all model families and model sizes. It’s worth noting that despite MiniLLM incorporating language modeling loss through the corpus used for pre-training, PromptKD exhibits better performance.
@@ 4.3 Analysis
Instruction Tell me whether these are books by black authors: I Know Why the Caged Bird Sings, Homegoing, Between the World and Me, Becoming, Beloved, The Color Purple.
Ground Truth The books: I Know Why the Caged Bird Sings, Homegoing, Between the World and Me, Becoming, Beloved, and The Color Purple are all written by black authors.
Teacher All of the books you mentioned are by black authors. I Know Why the Caged Bird
w/o Prompt Sings and Beloved are by Maya Angelou. Homegoing is by Yaa Gyasi. Between the World and Me is by Ta-Nehisi Coates. Becoming is by Michelle Obama. The Color Purple is by Alice Walker.
Teacher I Know Why the Caged Bird Sings, Homegoing, Between the World and Me, Be-
w/ Prompt coming, and The Color Purple are all books by black authors.
Student Yes, these are all books by black authors.
Table 2: Qualitative results of generated response from the Dolly validation set with and without using prompts for the Llama-13B teacher. A teacher with a prompt generates a response more similar to that of the student.
@@
SFT 15.70 28.90 15.70
KD 40.13 52.82 20.62
SeqKD 40.13 52.82 20.13
GKD 41.99 56.13 25.37
MiniLLM 68.91 78.54 85.71
PromptKD 43.62 56.57 26.97
Table 3: Comparison of computational costs. Where MA denotes the maximum allocated memory on the GPU and CA denotes the maximum cached memory on the GPU. Time indicates the total training time for each method. All computational costs are calculated on 4 NVIDIA A100 80 GB (PCIe) GPUs.
@@
To demonstrate the efficiency of PromptKD, we compare its computational cost with baselines in Table 3. OPT-13B and OPT-6.7B are used as the teacher and the student, with measurements conducted on 4 NVIDIA A100 80GB (PCIe) GPUs. From a time perspective, methods that sample the student at each iteration to create pseudo-targets take significantly more time than those that do not. In particular, MiniLLM requires a significant amount of time, primarily due to the additional use of the corpus used for pre-training, along with the complexity of calculating intricate rewards for optimization with policy gradient, unlike other methods. For the same reason, MiniLLM demands a substantial amount of memory. In contrast, PromptKD adds only a minimal amount of memory by introducing parameters equivalent to the product of prompt length and input embedding dimension. PromptKD demonstrates clear efficiency over MiniLLM and comparable costs to GKD, while significantly outperforming both in terms of performance. Therefore, PromptKD proves competitive in this regard.
Student-friendly knowledge
To provide a clear interpretation of student-friendly knowledge, we investigate how the prompt modifies the teacher model. As shown in Table 2, we generate responses to a validation set that was unseen during training using both teacher models—with and without prompt—and the trained student model. The findings reveal that while the original teacher generates a complex response, the student-friendly teacher, modified by the prompt, produces a response that is similar to and easily understood by the student. Notably, despite its simplicity, this response remains accurate. Furthermore, when modifying the teacher using the prompt, the quantitative verification of maintaining quality while achieving similarity to the student in responses is detailed in Appendix E. Therefore, the student-friendly knowledge distilled in PromptKD refers to knowledge transferred by a student-friendly teacher, who maintains a similar output distribution to the student for easier understanding while preserving the original generative performance. This aligns with the concept of adaptive teaching that served as the inspiration.
Ablation study
Due to the page limit, we detail an ablation study on regularization loss, prompt settings, and KL divergence in Appendix D.
@@ 5 Conclusions
In this work, we have pioneered the exploration of extracting and distilling student-friendly knowledge for generative language models. To achieve this, we have proposed a novel method called PromptKD, which leverages prompt tuning in knowledge distillation for the first time. Owing to the memory-efficient nature of prompts and the advantage of replacing full-parameter fine-tuning, particularly beneficial for larger models like LLMs, PromptKD has proven to be an efficient approach. Through extensive experiments, PromptKD has achieved state-of-the-art performance, confirming the effectiveness of student-friendly knowledge in generation tasks. Additionally, through exposure bias analysis, we have demonstrated that PromptKD successfully alleviates exposure bias throughout the training process.
@@ Appendix A Training Details
In our study, we employ the AdamW (Loshchilov and Hutter, 2019) optimizer for training, with batch sizes of 32 for GPT-2 Base and 8 for GPT-2 Medium and Large. The learning rates of prompt and student are set at 5e-5 for Base, 1e-5 for Medium, and 5e-6 for Large. In both the Llama and OPT model families, we set the batch size to 64 and the learning rates of prompt and student to 5e-6. For the generation, we sample with top-k and top-p parameters at 0 and 1.0, respectively, and use a fixed temperature of 1.0. Training and generation phases both have a maximum sequence length of 512 and a maximum prompt length of 256. Please note that we pre-process each instruction-following dataset by converting the instruction-response pairs into a standardized sentence structure, as shown in Table 4. For the reproducibility of our PromptKD, we will make both the code and the checkpoints public.
@@
Appendix B Qualitative Results
For the qualitative results, we present samples generated by student models trained using various methods. The samples are drawn from the S-NI dataset and utilize GPT-2 XL as the teacher model, with GPT-2 Large employed as the student model. Results are shown in Table 5. Additionally, the generation results obtained using the Llama model are summarized in Table 6.
Case 1
Instruction Convert a disfluent question to a proper question. A disfluent question is a question that has some interruptions in it while framing. A proper question is the correct form of the question without any disfluency.
Input What do no I meant Who do some tax agents want to make an impression on during an arrest?
Ground Truth Who do some tax agents want to make an impression on during an arrest?
SFT Some tax agents want to make an impression on during an arrest.
KD The tax agents want to make an impression on during an arrest.
SeqKD No I meant to ask who are some tax agents that want to make an impression on an arrest.
GKD No I mean that there are no tax agents who want to make an impression on the arrest.
MiniLLM A disfluent question is one that has some interruptions while framing. A proper questions is the correct form of a question without any disfluidity.
PromptKD Who do some tax agents want make an impression on during a arrest?
Case 2
Instruction Given an entity, a before event, an after event, and an attribute related to the entity, generate a sentence as output. Your sentence should show the changes in the attribute of the entity.
Input entity: rope, before: separate from ball, after: through slot in ball, attr: location
Ground Truth The location of the rope was separate from the ball before and through a slot in the ball afterwards.
SFT Rope was previously connected to a ball through a slot in the ball. When the rope is pulled through the ball, it changes location and is no longer connected to the ball.
KD Rope is a string. before: ball, after: rope stroke
SeqKD Rope is made of rope.
GKD before: separate from Ball, after: through slot In Ball, attr: location
MiniLLM through slot in ball
PromptKD Rope was separate from Ball until through slot in Ball.
Case 1
Instruction You are given two passages. For each word in the first passage, you need to find its synonym from the second passage such that both words have the same meaning and belong to the same part of speech. The input consists of two paragraphs with a list of words following each paragraph. The output is a list of synonyms (one for each word in the first paragraph) such that they have similar meaning as corresponding words in the first paragraph and belong to same parts of speech. Note: Some sentences may not contain a particular type of word
Input Paragraph 1: They water their crops with canal water. Paragraph 2: They irrigate their plants with pipeline fluid.
Ground Truth Canal water and pipeline fluid are both synonyms for irrigation.
MiniLLM Paragraph 2 synonyms for Paragraph 1: They irrigate Their plants with pipeline fluid.
PromptKD Canal water, pipeline fluid
Case 2
Instruction You are given a list of numerical values. Find the value that is halfway between the maximum and minimum values in the list, and output that value.
Input List = [1, 3, 4, 6, 8].
Ground Truth The value halfway between the maximum and minimum values in the list is 4.
MiniLLM The value that is halfwaybetween the maximum and minimum values is 4. This is because 4 is the average of the list, which is 1 + 3 + 4 + 6 + 8 / 5 = 4.
PromptKD The median value is 4.
Table 6: Sample response from the UnNI dataset. The Llama model is used here. Similar to the results with GPT-2, PromptKD generates appropriate responses following the instructions better than MiniLLM.
Appendix C GPT-4 Feedback Score
We follow the approach described in Appendix D.1 of MiniLLM (Gu et al., 2024) to measure the GPT-4 feedback score. We utilize the GPT-4 model with a temperature of 0.7. To evaluate model output compared to ground truth response, we employ a fixed form of prompt consisting of instruction, input, assistant 1, and assistant 2. The instruction of task and input are entered first, followed by the model output in assistant 1 and the ground truth response in assistant 2, as shown in Table 7. Through this prompt, scores for the model output and ground truth response, which are separated by spaces and range from 1 to 10, are obtained. The sum of the model output scores is divided by the sum of the ground truth scores to calculate the GPT-4 feedback score for each method. Similar to the main result in Table 1, scores are calculated for seeds 10, 20, 30, 40, and 50, then the average is taken. Following this approach, we measure the GPT-4 feedback scores for MiniLLM (Gu et al., 2024), GKD (Agarwal et al., 2024), and PromptKD, which demonstrated strong performance in Table 1. Here, we omit KD (Sanh et al., 2019) and SeqKD (Kim and Rush, 2016) from the measurement since they did not compete well against other baselines.
Prompt Format
### Instruction:
{instruction}
### Input:
{input}
### Assistant 1:
{model output}
### Assistant 2:
{ground truth response}
We would like to request your feedback on the performance of two AI assistants in response to the user instruction and input displayed above.
Please rate the helpfulness, relevance, accuracy, and level of detail of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space.
In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Table 7: Prompt format used for measuring GPT-4 feedback scores.
The evaluation results when using GPT-2 XL (1.5B) as the teacher and GPT-2 Large (760M) as the student are summarized in Table 8. Consistent with the trends observed in Table 1, PromptKD exhibits the best performance, followed by MiniLLM and then GKD. Particularly noteworthy is that PromptKD outperforms others on all datasets, further demonstrating the effectiveness of student-friendly knowledge.
Method Dolly SelfInst Vicuna
GKD 68.83 63.87 66.68
MiniLLM 71.39 66.96 67.78
PromptKD 72.12 67.22 68.01
Table 8: Evaluation results with GPT-4 feedback scores.
Appendix D Ablation Study
Regularization loss
#Params w/o $\mathcal{L}_{\text{reg}}$ w/ $\mathcal{L}_{\text{reg}}$
120M 21.97 22.25
340M 24.13 23.92
760M 24.47 25.08
Prompt settings
Although the regularization loss effectively mitigates the initial instability, the prompt’s length and initialization also significantly influence the prompt tuning process (Hou et al., 2022). Therefore, the average instruction-following performance is measured by varying the prompt length $m$ from 5, 7, 10 and the initialization method from random, padding, text. GPT-2 Large (760M) and GPT-2 XL (1.5B) are utilized for this ablation study. Results are summarized in Figure 4. In the padding method, all prompt tokens are initialized with the embedding of the "[PAD]" token, while in the text method, the sentence "Suppose you are a student." is tokenized, and these embeddings are used for initializing prompt tokens from the beginning. In this case, if the number of prompt tokens is smaller, the sentence is truncated, while if it is larger, all embeddings of the sentence are assigned, and then the embeddings are assigned again from the beginning for the next prompt token.
KL divergences